# Towards Vision Enhancing LLMs: Empowering Multimodal Knowledge Storage and Sharing in LLMs

### v1

@@ 4.5 Ablation Study and Analysis
Effects of MKS2. Comparing the experimental results of MKS2 w/o Multimodal-SFT, MKS2 w/o Multimodal-SFT & MoMEs, and Llama-2-7b-INST-LoRA in Table 1, we observed that the incorporation of visual information into the MVM positively impacted the model’s performance on common sense reasoning tasks. This demonstrates that MoMEs can leverage stored visual information to improve its understanding and reasoning abilities in various contexts. Additionally, the performances of MKS2 variants w/o Text-SFT and w/o Text-SFT + MoMEs on multimodal tasks further emphasizes its ability to access textual knowledge without significant compromise. The integration of visual information alongside textual data did not hinder the model’s capacity to extract and utilize textual knowledge effectively. This suggests that the core text-related capabilities of LLMs remain robust when dealing with multimodal inputs.
Impact of multimodal instruction data for MKS2 performance. Our experimental results in Table 1 highlights the impact of multimodal instruction data on MKS2’s performance. While it effectively enhances the accuracy in addressing questions related to physical world knowledge, it does not provide a substantial boost in solving intricate and complex problems that demand advanced reasoning and strategic thinking. These findings emphasize the importance of tailoring data and approaches to specific task requirements when leveraging multimodal data in large language models for optimal performance. Further research may uncover strategies to bridge this gap for complex problem-solving tasks.
Table 3: Ablation Experiments on five datasets requiring common sense. We explore various data and visual memory sizes to check model performances. All models are built upon Llama-2-7b, where ${}^{\clubsuit}$ signifies MKS2 w/o Multimodal-SFT. ”-AM-BM” indicates that it includes a ”A” size of visual memory and was trained using ”B” number of image-text pairs.
@@
${}^{\dagger}$ 31.62 35.3 64.90 23.53 42.37
${}^{\dagger}$ 42.58 38.2 53.42 29.32 39.71
${}^{\clubsuit}$-410M-2.3M 44.06 50.5 67.10 28.99 57.72
${}^{\clubsuit}$-410M-12.3M 46.12 50.8 64.68$\downarrow$ 30.06 55.31$\downarrow$
${}^{\clubsuit}$-810M-2.3M 43.78 48.6 66.97 28.89 57.21
${}^{\clubsuit}$-810M-12.3M 46.56 49.8 68.10 30.65 57.84
Figure 4: An illustration of cases generated by comparative models and MKS2. Green words refer to the correct answer in the natural question answering tasks. Yellow parts represent the interesting and correct content in the response.
Could text-only instruction data be effective for enhancing multimodal performance while building MLLMs? In the ablation experiments discussed in Table 2, text-only instruction data was found to enhance the performance of pre-trained LLMs on various open multimodal problems, particularly open-ended questions that demand a fusion of textual and visual information. However, the introduction of text-only instruction data may lead to a trade-off, as it could potentially diminish the LLMs to answer visual questions involving scene text recognition. These findings underscore the importance of striking a careful balance when incorporating additional data modalities into LLMs, with a keen consideration of task requirements and data characteristics. Such careful integration ensures that overall model performance is optimized without unintended consequences on its ability to handle diverse multimodal challenges. Consequently, optimizing LLMs for multimodal tasks requires a nuanced approach tailored to the specific problem domain.
@@
Impact of visual memory size and data scale. We introduce more image-text pairs from LAION-400M (Schuhmann et al., 2021) to analyse the impacts of vision memory and data sizes, and the experimental results are shown in Table 3. Our experimental results reveal that enlarging the size of the pre-trained image-text data leads to improvements in the MKS2 model’s performance across various downstream tasks. Additionally, we observe that expanding the visual storage module of the MKS2 model not only enhances performance but also contributes to greater stability in this enhancement, as the size of the image-text data increases. To summarize, these findings suggest a twofold strategy for optimizing the MKS2-based LLMs during the supervised fine-tuning phase: increasing the size of the image-text data and selecting a proportionately larger visual storage size. This approach is crucial for achieving more consistent and stable performance improvements.
@@ 4.6 Case Study
We present some cases in Figure 4 to further show the overall capability of MKS2-Llama-2. We can see that the proposed model achieve better performance while answering commonsense Q&A with physical knowledge. In addition, we also observe that the multimodal understanding capability of MKS2-Llama-2 is powerful, such as it could recognize the funny of the cat with a lion’s head. In addition, it can employ the relevant knowledge to enrich the response based on visual clues, e.g., the short story around the content shown in the second image.
5 Related Works
Visual knowledge enhanced methods. There is a long line of work on utilizing explicit visual information to improve the imaginative representation of language, thus promoting diverse generation capability of LLMs. Particularly, Jin et al. (2022) leverage visual knowledge in NLP tasks, developing multiple cross-model enhanced methods to improve the representation capability of pretrained language models. Some works (Shi et al., 2019; Lu et al., 2022; Li et al., 2023b) proposed to retrieve images corresponding to texts from the image corpus and use visual knowledge to improve the performance on the downstream tasks such as text completion (Zellers et al., 2019), story generation (Fan et al., 2018), and concept-to-text (Barzilay & Lapata, 2005). Recently, some researchers (Long et al., 2021; Yang et al., 2021; Zhu et al., 2022) proposed to utilize the powerful text-to-image technical to obtain the imagination representation of language and infuse them into the language model via the prefix-tuning way. In this paper, we present visual information storage in LLMs and achieve visual knowledge enhancing LLMs without explicitly inputting images to language models.
LLMs for vision. Recent works (Zhang et al., 2023; Zhu et al., 2023; Li et al., 2023a) towards multimodal LLMs focus on utilizing the extensive knowledge and language generation capabilities of LLMs to solve multimodal tasks (Li et al., 2023d), especially for visual understanding and reasoning. Firstly, these works usually map the visual information obtained by pretrained visual encoder into the representation space of LLMs, through a learnable linear projection layer (Merullo et al., 2023), MLP, or Q-Former (Li et al., 2023a). This stage is usually called feature alignment and only a few hundred thousand data may be needed to do a good job (Liu et al., 2023a). Afterwards, the initial MLLMs will be tuned via multimodal instruction-following data (Ye et al., 2023; Li et al., 2023c; Bai et al., 2023). At this stage, LLMs and projection layer are often tuned together only with multimodal instruction data. The commonly used large language model is the SFT LLM and the tuning approach adopts widely used lightweight LoRA (Hu et al., 2021). These works, however, rarely consider using visual knowledge to enhance the pure text processing capabilities of LLMs, thereby building a more robust LLM or MLLM.
@@ 6 Conclusion
In this paper, we present a new approach MKS2 that allows LLMs to memorize and employ visual information, achieving multimodal knowledge storage and collaboration in LLMs. MKS2 consists of modular visual memory and soft mixtures-of-multimodal experts, which are used to store visual information and realize multimodal knowledge collaboration, respectively. We conduct extensive experiments on many NLP and VQA tasks and the experimental results show that MKS2 is capable of enhancing the reasoning capability of LLMs and being used to solve multimodal problems.

### v2

@@ 6.6 Ablation Study and Analysis
Effects of MKS2. Comparing the experimental results of MKS2 w/o Multimodal-SFT, MKS2 w/o (Multimodal-SFT & MoMEs), and Llama-2-7b-INST-LoRA in Table 1, we observed that the incorporation of visual information into the MVM positively impacted the model’s performance on common sense reasoning tasks. This demonstrates that MoMEs can leverage visual knowledge from MVM to improve their understanding and reasoning abilities in various contexts. As shown in Table 1, we performed ablation studies using both the 7B and 13B versions of the Llama model, comparing settings with and without the MVM. For instance, the condition “without Multimodal-SFT” was compared to “without (Multimodal-SFT & MoMEs)“. In the absence of the MVM module, the model loses its mixture-of-experts capability and reverts to standard fine-tuning, trained solely on plain text with a rank $r=$ 8 configuration, e.g., w/o (Multimodal-SFT & MoMEs). The comparison between these two variants demonstrates that introducing MVM-which stores visual information as part of a visual expert-leads to noticeable improvements on textual tasks. For example, we observe an average performance gain of 6% on StrategyQA. This suggests that visual knowledge stored via MVM provides complementary information that aids in textual reasoning. Additionally, the performances of MKS2 variants w/o Text-SFT and Text-SFT & MoMEs on multimodal tasks further emphasize its ability to access textual knowledge without significant compromise. The integration of visual information alongside textual data did not hinder the model’s capacity to extract and utilize textual knowledge effectively. This suggests that the core text-related capabilities of LLMs remain robust when dealing with multimodal inputs.
Table 5: Model performances on MMBench. The language Model and Vision Model represent the backbones of MLLMs. “TotalParams” indicates the total parameters of MLLMs. Logical Reasoning (LR), Attribute Reasoning (AR), Relation Reasoning ($R$), Fine-grained Perception (Cross Instance) (FP-C), Fine-grained Perception (Single Instance) (FP-S), and Coarse Perception (CP).
@@
OpenFlamingo (Alayrac et al., 2022) - - 9B 4.3 9.1 11.4 3.3 2.5 1.6 1.5
OpenFlamingo v2 (Alayrac et al., 2022) - - 9B 5.7 11.4 12.8 1.4 5.5 0.8 4.0
MiniGPT-4 (Zhu et al., 2024) Vicuna 7B EVA-G 1.1B 8B 23.0 13.6 32.9 8.9 28.8 11.2 28.3
MMGPT (Gong et al., 2023) LLaMA 7B CLIP VIT-L/14 428M 9B 16.0 1.1 23.8 20.7 18.3 5.2 18.3
PandaGPT (Su et al., 2023) Vicuna 13B ImageBind ViT-H/14 1.2B 14B 30.6 15.3 41.5 22.0 20.3 20.4 47.9
VisualGLM (Du et al., 2022) ChatGLM 6B EVA-G 1.1B 8B 33.5 11.4 48.8 27.7 35.8 17.6 41.5
InstructBLIP (Dai et al., 2024) Vicuna 7B EVA-G 1.1B 8B 33.9 21.6 47.4 22.5 33.0 24.4 41.1
LLaVA (Liu et al., 2023a) LLaMA 7B CLIP VIT-L/14 428M 8B 36.2 15.9 53.6 28.6 41.8 20.0 40.4
LLaMA-Adapter-v2 (Gao et al., 2023) LLaMA 7B CLIP VIT-L/14 428M 8B 38.9 7.4 45.3 19.2 45.0 32.0 54.0
G2PT LLaMA 7B ViT-G 1.1B 8B 39.8 14.8 46.7 31.5 41.8 34.4 49.8
mPLUG-Owl (Ye et al., 2023) LLaMA 7B CLIP VIT-L/14 428M 8B 56.7 30.7 65.7 52.1 61.0 45.6 65.1
Otter-I (Li et al., 2023a) LLaMA 7B CLIP VIT-L/14 428M 9B 48.3 22.2 63.3 39.4 46.8 36.4 60.6
Shikra (Chen et al., 2023) Vicuna 7B CLIP VIT-L/14 428M 8B 60.2 33.5 69.6 53.1 61.8 50.4 71.7
LMEye (Li et al., 2024a) Vicuna 7B CLIP VIT-L/14 428M 8B 61.0 37.5 70.8 43.6 62.1 55.9 73.2
LLaVA-v1.5 (Liu et al., 2024a) Vicuna 7B CLIP VIT-L/14 428M 8B 62.5 22.9 72.4 53.9 63.8 58.7 75.7
MKS2-Llama-2-7b ($r=16)$ Vicuna 7B CLIP VIT-L/14 428M 8B 73.0 57.1 72.7 55.6 80.0 75.8 75.7
@@
Impact of visual memory size and data scale. We introduce more image-text pairs from LAION-400M (Schuhmann et al., 2021) to analyze the impacts of vision memory and data sizes, and the experimental results are shown in Table 4. Our experimental results reveal that enlarging the size of the pre-trained image-text data leads to improvements in the MKS2 model’s performance across various downstream tasks. Additionally, we observe that expanding the visual storage module of the MKS2 model not only enhances performance but also contributes to greater stability in this enhancement, as the size of the image-text data increases. These findings suggest a twofold strategy for optimizing the MKS2-based LLMs during the SFT phase: increasing the size of the image-text data and selecting a proportionately larger visual storage size. It is crucial for achieving consistent and stable performance improvements.
Could text-only instruction data be effective for enhancing multimodal performance while building MLLMs? As discussed in Table 2, text-only instruction data was found to enhance the performance of pre-trained LLMs on various open multimodal problems, particularly open-ended questions that demand a fusion of textual and visual information. However, the introduction of text-only instruction data may lead to a trade-off, as it could potentially diminish the LLMs to answer visual questions involving scene text recognition. These findings underscore the importance of striking a careful balance when incorporating additional data modalities into LLMs, with a keen consideration of task requirements and data characteristics. Such careful integration ensures that overall model performance is optimized without unintended consequences on its ability to handle diverse multimodal challenges.
Impact of multimodal instruction data. Our experimental results in Table 1 highlight the impact of multimodal instruction data on MKS2’s performance. While it effectively enhances the accuracy in addressing questions related to physical world knowledge, it does not provide a substantial boost in solving intricate and complex problems that demand advanced reasoning and strategic thinking. These findings emphasise the importance of tailoring data and approaches to specific task requirements when leveraging multimodal data in LLMs for optimal performance. This may also be because the current visual instruction data is mainly oriented towards visual understanding, and there is less data on visual background knowledge and reasoning. We acknowledge that differences in pre-training data scale and quality may affect model performance. It is worth noting that models such as LLaVA (0.6M images), LLaVA-v1.5 (1M images), and BLIP-2 (123M images) indeed use varying amounts of pre-training data. Nevertheless, LLaVA achieve better performance than BLIP-2 after being fine-tuned on high-quality multimodal instruction data, underscoring the significant role that instruction tuning plays in the downstream task. The primary goal of pre-training in multimodal LLMs is to align visual representations with the semantic space of the language model, thereby enabling better cross-modal understanding. Hence, to ensure a fair comparison with LLaVA-V1.5, we used the same instruction tuning dataset during fine-tuning. This allows us to more directly evaluate the effectiveness of our architectural improvements, rather than reflecting disparities in pre-training data. In Table 2, our final model MKS2-Llama-2-7b achieves better performance with the same scale of parameters and data compared to LLaVA-v1.5. It is worth noting that utilizing only a minor segment of visual instruction tuning data in the SFT stage of MKS2 will not effectively enhance the model’s multimodal performance.
Table 6: Performance metrics across LoRA with different ranks ($r=$). FLOPs indicates the forward propagation latency FLOPs per GPU. Params Count refers to the total tuning parameters and the proportion of total model parameters.
Rank ($r=$) 8 16 32 64
FLOPs 61.0T 62.8T 67.0T 66.3T
Training Time 23h 23h 23h 24h
Params Count 17.8M/0.25% 35.7M/0.50% 71.5M/1.00% 143.3M/2.00%
MMBench 70.0 73.0 71.9 69.3
Figure 6: An illustration of cases generated by comparative models and MKS2. Green words refer to the correct answer in the natural question-answering tasks. Yellow parts represent the interesting and correct content in the response.
MoEs for multimodal understanding and reasoning. We also explored the MoE architecture-based model, MKS2-Llama-2-7b, focusing on its multimodal understanding and performance. Our model utilizes the same pretraining and visual instruction tuning data as LLaVA-V1.5, with the key distinction being the integration of a visual memory module during the pretraining phase. As shown in Tables 5 and 2, our MoE-based multimodal storage and sharing approach outperforms the dense language model alone when constructing a multimodal large model. Additionally, our model did not demonstrate significant improvements with a limited set of visual instruction data in Table 2; however, when a larger multimodal dataset LLaVA-665k was used, we observed substantial enhancements in performance on the inference dataset MMBench (Table 5) and traditional VQA tasks, such as VQAv2 and OK-VQA. This suggests that activating the collaborative capabilities of multimodal experts necessitates more data. The overall training time was about 12 hours on an 8-card A100-80G setup, which may be a worthwhile investment for developing more robust MLLM.
Model performance with different LoRA rank ($r$). We conducted an empirical study to investigate the relationship between the number of learnable parameters in LoRA and model performance, which is shown in Table 6. Our analysis of LoRA parameters reveals two key findings: 1) Performance Saturates: Model performance on MMBench does not improve significantly with more parameters.
Figure 7: An illustration of cases. Words with a green background refer to the correct answer, yet Purple words are results inferred by MKS2.
At a high rank (r>r>16), performance degrades, suggesting excessive parameters lead to overfitting or optimization issues for this task. 2) Efficient Computation: For moderate ranks (r=r=8 to 32), increased parameters did not slow training. The added computational load appears to better utilize GPU capacity, maintaining throughput until the parameter count becomes excessive. The overall training time of the model is of the same order of magnitude, and the model performance increase in downstream tasks is not obvious and achieves best on $r=16$. Hence, this suggests that warming up expert activations with limited data does not require a large number of learnable parameters.
@@ 6.7 Case Study
We present some cases in Figure 6 to further show the overall capability of MKS2-Llama-2 in multimodal understanding and text-only question answering. We can see that the proposed model achieves better performance while answering commonsense Q&A with physical knowledge, as the above examples shown in Figure 6. In addition, we also observe that the multimodal understanding capability of MKS2-Llama-2 is also powerful, such as recognising the fun of the cat with a lion’s head in the third image. It could employ the relevant knowledge to enrich the response based on visual clues, e.g., the short story around the content shown in the second image with a young man and a woman. These findings are consistent with the experimental results in the above tables.
However, as shown in Figure 7, we find that our model still struggles with some knowledge-based visual questions, where these questions require the model to combine its visual understanding with external, factual knowledge that is not explicitly contained in the image pixels, and visual-spatial reasoning questions, which demands model to understand and reason about the spatial relationships, orientations, or relative positions of objects within an image. Based on this finding, the model’s primary limitations are its deficiencies in knowledge-based visual reasoning and complex visual-spatial reasoning. While it succeeds at straightforward visual recognition (e.g., identifying a coastal landscape), it fails to integrate external world knowledge, as shown by its inability to apply basic geographical concepts to a map. Furthermore, it struggles to interpret dynamic interactions and spatial relationships between objects, erroneously identifying a clear predator-prey scenario as a symbiotic relationship, highlighting a weakness in inferring actions and causal narratives from visual scenes.
@@ 7 Conclusion
In this paper, we present a new approach named MKS2, that allows LLMs to memorize and employ visual information, achieving multimodal knowledge storage and collaboration in LLMs. MKS2 consists of modular visual memory and a soft mixture of multimodal experts, which are used to store visual information and realise multimodal knowledge collaboration, respectively. To make the visual memory effective, we introduce the language-centric learning framework that includes text-image understanding and generation. We conduct extensive experiments on many natural language processing and visual question-answering tasks. The experimental results show that MKS2 is capable of enhancing the reasoning capability of LLMs and can be used to solve multimodal problems. The ablation study suggests that introducing the MoE architecture in dense pretrained LLMs could improve their language understanding and multimodal performances.
8 Acknowledgment
This work is jointly supported by National Natural Science Foundation of China (Grant No. 62422603), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024B0101050003), and Shenzhen Science and Technology Program (Grant No. ZDSYS20230626091203008).
9 Contributors
Contributors
Yunxin Li, Zhenyu Liu, Baotian Hu, Yuxin Ding, and Min Zhang are with the School of Computer Science and Technology, Harbin Institute of Technology, Shenzhen. (e-mail: liyx@hit.edu.cn, liuzhenyuhit@gmail.com, hubaotian@hit.edu.cn, yxding@hit.edu.cn, and zhangmin2021@hit.edu.cn).
Wei Wang and Xiaochun Cao are with the School of Cyber Science and Technology, Shenzhen Campus of Sun Yat-sen University. (e-mail: wangwei29@mail.sysu.edu.cn and caoxiaochun@mail.sysu.edu.cn)
Corresponding Author
Baotian Hu
Harbin Institute of Technology, Shenzhen
Email: hubaotian@hit.edu.cn